[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jun 23, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jun 23 19:00:00 UTC 2017


HSU Shift Log for Fri Jun 23 2017 14:59:54 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Rebooted tilefish since it had been up for 77 days. It  
didn't come up properly after reboot, so I put in a help desk ticket and  
Doug from NCO resolved the issue quickly.

AWIPS-2/NCP issues:

Other Techincal issues: The SHIPS guidance for EP93 wasn't able to find  
satellite data for 12Z or 18Z, so its using proxy data.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170623/52760af4/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list